# Forecasting future anatomies: Longitudinal brain MRI-to-MRI Prediction

### v1

@@ 1 Introduction
Understanding how the human brain changes over time is a central goal in neuroimaging research. Longitudinal magnetic resonance imaging (MRI) provides rich insights into brain development, aging, and structural alterations associated with neurological and psychiatric conditions. In addition, predicting future MRI patterns is valuable because anticipated brain changes can serve as early indicators of disease progression or cognitive decline, offering a data-driven pathway toward forecasting future clinical outcomes. However, longitudinal datasets are often limited in temporal coverage and sample size, constraining our ability to model individual trajectories of brain change. To address these challenges, recent research has begun exploring computational approaches that can generate future brain MRI scans, effectively trying to predict how an individual’s brain structure will evolve over time [1, 2]. Accurate prediction of future brain morphology could open new avenues for studying brain aging and disease progression, improving patient-level prognosis.
@@
Traditionally, machine learning in neuroimaging has focused on predicting diagnostic [3] or cognitive outcomes from imaging features [4, 5]. Although valuable for clinical applications, these approaches do not offer predictions on how brain anatomy itself changes over time. In contrast, image-to-image prediction seeks to model the full spatial evolution of brain structure, learning how an individual’s MRI transforms over time. This task requires capturing both spatial anatomy and temporal dynamics, making it a challenging and highly informative problem in computational neuroimaging.
@@
A rapidly expanding body of research has explored deep learning approaches for medical image-to-image translation. In neuroimaging, this work has focused primarily on cross-modality synthesis, for example, generating a pseudo-FDG positron emission tomography (PET) image from an MRI or vice versa [6, 7]. Many MRI synthesis methods using deep learning translate between MRI pulse sequences [8], most commonly T1wT2w/FLAIR, showing that one sequence can often predict another’s contrast. However, they do not address longitudinal prediction within the same modality. Another related line of research involves MRI-to-MRI image synthesis, in which generative models simulate brain scans under different conditions, such as aging or disease effects, but do not predict the actual future MRI of each individual. Ravi et al. introduced degenerative adversarial neuroimage nets to produce aged brain images constrained by disease patterns [2], and Xia et al. (2021) proposed an autoencoder-based method to generate aged MRIs from a single baseline scan [1]. Although these models generate visually convincing progressions, they typically capture average or prototypical changes rather than individualized, subject-specific evolution. More recently, Pombo et al. (2023) combined conditional generative models with diffeomorphic transformations to create “counterfactual” MRIs under altered conditions [9, 10]; however, such approaches are not trained to reproduce each participant’s true future brain image.
@@
In this study, we investigate longitudinal MRI-to-MRI prediction, enabling voxel-level modeling of individual brain changes over time. Unlike prior work focused on regional measures [11] or population-average trajectories[1], our approach predicts the entire future MRI of each individual, capturing the complex spatial pattern of anatomical evolution. In particular, we implement and compare five deep learning architectures within a unified framework, evaluate their predictions against follow-up scans using both global and regional metrics, and test their generalizability on an independent cohort.
@@ 2.1 Datasets and Preprocessing
Data used in this work were obtained from the ADNI (http://adni.loni.usc.edu) and the AIBL study group (https://aibl.org.au). For up-to-date information, see (www.adni-info.org). AIBL study methodology has been reported previously [12].
We used longitudinal T1-weighted MRI from ADNI and AIBL, including data from all the participants who had undergone at least two scans. Images were preprocessed using CAT12 [13] with segmentation, spatial normalization, modulation, smoothing ( FWHM), and downsampling to , yielding gray matter density maps. Two ADNI training sets were generated: BigDataset (2746 pairs from 1255 participants with 24-month intervals at varying starting timepoints) and SmallDataset (1191 baseline-to-month-24 pairs). Data were split by participant (40 validation, 140 test) to prevent leakage. The ADNI test set consisted of baseline-to month-24-pairs. AIBL served as external validation with 18-month scans extrapolated to 24 months (173 participants).
@@ 2.2 Deep Learning Architectures, Training and Evaluation
We implemented five architectures predicting 3D MRI at from : UNet [14] with encoder-decoder and skip connections; U2-Net [15] with nested residual U-blocks for multi-scale features; UNETR [16] using Vision Transformer encoder with convolutional decoder; Time-Embedding UNet (TEUNet)[17] with learnable time embeddings; ODE-UNet [18, 19] models continuous latent dynamics via ; integrating this neural ODE from baseline to months yields the predicted future scan. Here is the latent state, is time (months), and is the learned vector field.
UNet, U2-Net, UNETR, and TEUNet used Mean Squared Error (MSE) loss. ODE-UNet employed following ImageFlowNet [17], where is a perceptual/feature-matching term (using a frozen encoder) to preserve anatomy/texture, penalizes deviations from the latent ODE along the 24-month trajectory, and are scalar weights set to in our experiments. All models used the Adam optimizer (learning rate , weight decay , batch size 1 3D MRI) with batch normalization and early stopping on A100 GPUs. Each model was trained with a fixed maximum training time of 36 hours. Evaluation metrics were MSE, Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM). To study how well the models predict changes, we used -Pearson correlation for longitudinal change. Define predicted image by and vectorized in-mask change maps and . Then participant-wise (Global) -Pearson correlation is the average Pearson correlation .

### v2

@@ 1 Introduction
Understanding how the human brain changes over time is a central goal in neuroimaging research. Longitudinal magnetic resonance imaging (MRI) provides rich insights into brain development, aging, and structural alterations associated with neurological and psychiatric conditions. In addition, predicting future MRI patterns is valuable because anticipated brain changes can serve as early indicators of disease progression or cognitive decline, offering a data-driven pathway toward forecasting future clinical outcomes. However, longitudinal datasets are often limited in temporal coverage and sample size, constraining our ability to model individual trajectories of brain change. To address these challenges, recent research has begun exploring computational approaches that can generate future brain MRI scans, effectively trying to predict how an individual’s brain structure will evolve over time [xia2021learning, ravi2022degenerative]. Accurate prediction of future brain morphology could open new avenues for studying brain aging and disease progression, improving patient-level prognosis.
@@
Traditionally, machine learning in neuroimaging has focused on predicting diagnostic [spasov2019parameter] or cognitive outcomes from imaging features [moradi2025integratingnba, colliot2023machine]. Although valuable for clinical applications, these approaches do not offer predictions on how brain anatomy itself changes over time. In contrast, image-to-image prediction seeks to model the full spatial evolution of brain structure, learning how an individual’s MRI transforms over time. This task requires capturing both spatial anatomy and temporal dynamics, making it a challenging and highly informative problem in computational neuroimaging.
@@
A rapidly expanding body of research has explored deep learning approaches for medical image-to-image translation. In neuroimaging, this work has focused primarily on cross-modality synthesis, for example, generating a pseudo-FDG positron emission tomography (PET) image from an MRI or vice versa [dayarathna2024deep, sikka2025mri]. Many MRI synthesis methods using deep learning translate between MRI pulse sequences [dar2019image], most commonly T1wT2w/FLAIR, showing that one sequence can often predict another’s contrast. However, they do not address longitudinal prediction within the same modality. Another related line of research involves MRI-to-MRI image synthesis, in which generative models simulate brain scans under different conditions, such as aging or disease effects, but do not predict the actual future MRI of each individual. Ravi et al. introduced degenerative adversarial neuroimage nets to produce aged brain images constrained by disease patterns [ravi2022degenerative], and Xia et al. (2021) proposed an autoencoder-based method to generate aged MRIs from a single baseline scan [xia2021learning]. Although these models generate visually convincing progressions, they typically capture average or prototypical changes rather than individualized, subject-specific evolution. More recently, Pombo et al. (2023) combined conditional generative models with diffeomorphic transformations to create “counterfactual” MRIs under altered conditions [pombo2023equitable, fu2025synthesizing]; however, such approaches are not trained to reproduce each participant’s true future brain image.
@@
In this study, we investigate longitudinal MRI-to-MRI prediction, enabling voxel-level modeling of individual brain changes over time. Unlike prior work focused on regional measures [hadji2025predicting] or population-average trajectories[xia2021learning], our approach predicts the entire future MRI of each individual, capturing the complex spatial pattern of anatomical evolution. In particular, we implement and compare five deep learning architectures within a unified framework, evaluate their predictions against follow-up scans using both global and regional metrics, and test their generalizability on an independent cohort.
@@ 2.1 Datasets and Preprocessing
Data used in this work were obtained from the ADNI (http://adni.loni.usc.edu) and the AIBL study group (https://aibl.org.au). For up-to-date information, see (www.adni-info.org). AIBL study methodology has been reported previously [ellis2009australian].
We used longitudinal T1-weighted MRI from ADNI and AIBL, including data from all the participants who had undergone at least two scans. Images were preprocessed using CAT12 [gaser2024cat] with segmentation, spatial normalization, modulation, smoothing ( FWHM), and downsampling to , yielding gray matter density maps. Two ADNI training sets were generated: BigDataset (2746 pairs from 1255 participants with 24-month intervals at varying starting timepoints) and SmallDataset (1191 baseline-to-month-24 pairs). Data were split by participant (40 validation, 140 test) to prevent leakage. The ADNI test set consisted of baseline-to month-24-pairs. AIBL served as external validation with 18-month scans extrapolated to 24 months (173 participants).
@@ 2.2 Deep Learning Architectures, Training and Evaluation
We implemented five architectures predicting 3D MRI at from : UNet [ronneberger2015u] with encoder-decoder and skip connections; U2-Net [qin2020u2] with nested residual U-blocks for multi-scale features; UNETR [hatamizadeh2022unetr] using Vision Transformer encoder with convolutional decoder; Time-Embedding UNet (TEUNet)[liu2025imageflownet] with learnable time embeddings; ODE-UNet [chen2018neural, bilovs2021neural] models continuous latent dynamics via ; integrating this neural ODE from baseline to months yields the predicted future scan. Here is the latent state, is time (months), and is the learned vector field.
UNet, U2-Net, UNETR, and TEUNet used Mean Squared Error (MSE) loss. ODE-UNet employed following ImageFlowNet [liu2025imageflownet], where is a perceptual/feature-matching term (using a frozen encoder) to preserve anatomy/texture, penalizes deviations from the latent ODE along the 24-month trajectory, and are scalar weights set to in our experiments. All models used the Adam optimizer (learning rate , weight decay , batch size 1 3D MRI) with batch normalization and early stopping on A100 GPUs. Each model was trained with a fixed maximum training time of 36 hours. Evaluation metrics were MSE, Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM). To study how well the models predict changes, we used -Pearson correlation for longitudinal change. Define predicted image by and vectorized in-mask change maps and . Then participant-wise (Global) -Pearson correlation is the average Pearson correlation .